OpenStudy (anonymous):

In a study of 250 adults, the mean heart rate was 70 beats per minute. Assume the population of heart rates is known to be approximately normal with a standard deviation of 12 beats per minute. What is the 99% confidence interval for the mean beats per minute? 68.9 – 76.3 70 – 72 61.2 – 72.8 68 – 72

OpenStudy (anonymous):

I got it.

OpenStudy (kirbykirby):

For a 99% C.I., for have \(\alpha=0.01\). \[ \left[\bar{x}-Z_{\alpha/2}\frac{s}{\sqrt{n}}, \,\,\bar{x}+Z_{\alpha/2}\frac{s}{\sqrt{n}} \right]\] You can verify on a table that the corresponding z-value should be 2.575 \[ \left[70-2.575\cdot\frac{12}{\sqrt{250}}, \,\,70+2.575\frac{12}{\sqrt{250}} \right]\]

OpenStudy (anonymous):

@brooklynransome found the answer?